OpenStudy (anonymous):

Find the first four terms of the sequence represented by the expression. 3n + 5 Answer choices: a. 5, 8, 11, 14 b. 8, 11, 14, 17 c. 3, 6, 9, 12 d. 0, 8, 11, 14 Have no clue how to do this... PLZ PLZ HELP! thx! xx

OpenStudy (anonymous):

a

OpenStudy (anonymous):

How???

OpenStudy (ajprincess):

plug in 0,1,2,3 in place of n and find the value of 3n+5 for each of those values of n. Does that help?

OpenStudy (anonymous):

Yes that does help! So like... 3*0 = 0 + 5 = 5. Therefore the first answer would have to be 5? I think I got it!

OpenStudy (anonymous):

for the first term you put ( 0 ) instead of n , then for the second one put (1 ) instead of n and so on

OpenStudy (ajprincess):

ya that's right:)